### v1

# Non-linear traces of Choquet type on AF algebras

Abstract.
We study non-linear traces of Choquet type on AF algebras. Building on the characterization of Choquet traces on matrix algebras due to Nagisa–Watatani, we generalize the construction to arbitrary unital AF algebras. We show that there is a one-to-one correspondence between such traces and increasing functions on the dimension scale, and we obtain explicit Choquet formulas in terms of the spectrum and ranks of spectral projections along a fixed AF filtration.
1. Introduction
The Choquet integral provides a flexible extension of the Lebesgue integral in which the underlying set function is allowed to be non-additive but monotone. In the finite-dimensional discrete setting, a capacity $\mu$ on a finite set $\Omega$ gives rise to a map on $0,\infty)^{\Omega}$ which is characterized by three basic properties: monotonicity, positive homogeneity, and comonotonic additivity. This viewpoint has been extensively developed in the theory of non-additive measure and integral (see, for example, [4], [5], [7]), and has found applications in decision theory, risk analysis, and probability under ambiguity.
In the operator algebraic setting, it is natural to ask whether one can develop analogous “Choquet-type” integration theories for positive elements of a C-algebra. A first step in this direction was taken by Nagisa and Watatani [11], who introduced non-linear traces on full matrix algebras and studied them from the viewpoint of matrix analysis and majorization. In their framework, a non-linear trace on $M_{n}(\mathbb{C})$ can be characterized as a Choquet map $\varphi:M_{n}(\mathbb{C})^{+}\to[0,\infty)$ which is unitarily invariant, monotone, positively homogeneous, and comonotonically additive with respect to the spectral functional calculus. They proved that such maps admit explicit expressions in terms of the eigenvalues of a positive matrix and the ranks of its spectral projections, and that there is a one-to-one correspondence between the maps and the increasing functions on the finite dimension scale $\{0,1,\dots,n\}$.
Their results show that non-linear traces provide a natural non-commutative extension of discrete Choquet integrals to the matrix setting. On the one hand, every such trace is completely determined by its values on projections, and these values are encoded by an increasing function on the scale. On the other hand, given an increasing function on the scale, one obtains a non-linear trace via an explicit Choquet formula which involves eigenvalue gaps and the ranks of the corresponding spectral projections. In particular, ordinary tracial states appear as the special case where the scale function is linear, that is, $\alpha(k)=k\alpha(1)$ for any $k\in\{0,1,\cdots,n\}$.
Beyond the matrix case, Nagisa and Watatani have further extended the theory of non-linear traces of Choquet type to broader operator-algebraic settings. In particular, they studied non-linear traces on the algebras of compact operators and on semi-finite factors, where eigenvalues are replaced by generalized singular values and spectral distribution functions [12, 13]. More recently, they investigated probabilistic aspects of such traces, including a law of large numbers for non-linear traces of Choquet type on finite factors [14]. We refer the reader to these works for further generalizations of non-linear traces beyond the finite-dimensional framework considered in this paper.
We also note that several ideas in the present paper are inspired by Sogame’s master’s thesis[15]. In particular, the viewpoint of parametrizing such traces by increasing maps on the dimension scale and some arguments in our AF-extension are adapted from that thesis.
The aim of this paper is to extend this picture from full matrix algebras to arbitrary unital AF algebras. For an AF algebra $A$, the ordered $K_{0}$-group and its distinguished scale $\Gamma(A)$ play the role of a global dimension scale which encodes the sizes of projections along an AF filtration.
From a different viewpoint, the present work may also be regarded as a non-linear extension of the correspondence between traces on AF algebras and order-preserving states on the ordered $K_{0}$-group. In the linear setting, it is well known that tracial states on an AF algebra are described in terms of the ordered $K_{0}$-group and its scale, see for example Davidson’s book [6] (see also [8], [10]) and the work of Blackadar on traces on simple AF $C^{*}$-algebras [2]. Motivated by this perspective, we regard $\Gamma(A)$ as the natural parameter space for non-linear traces of Choquet type on a unital AF algebra.
More precisely, we consider maps $\varphi:A^{+}\to[0,\infty)$ satisfying the following properties:
  • unitary invariance: $\varphi(uau^{*})=\varphi(a)$ for all $a\in A^{+}$ and all unitaries $u\in A$;
  • monotonicity: $0\leq a\leq b$ implies $\varphi(a)\leq\varphi(b)$;
  • positive homogeneity: $\varphi(ta)=t\,\varphi(a)$ for all $t\geq 0$ and $a\in A^{+}$;
  • comonotonic additivity on the spectrum: for each $a\in A^{+}$ and comonotonic functions $f,g\in C(\sigma(a))^{+}$, we have $\varphi(f(a)+g(a))=\varphi(f(a))+\varphi(g(a))$.
We call such maps Choquet traces on $A$. When $A=M_{n}(\mathbb{C})$, this notion reduces to the non-linear traces of Nagisa–Watatani [11]. In the AF case, the extra structure coming from the inductive limit decomposition allows one to express Choquet traces in terms of spectral data inside each finite-dimensional building block and an increasing function on the global dimension scale.
The main result of this paper shows that there is a bijective correspondence between Choquet traces on a unital AF algebra $A$ and increasing functions on $\Gamma(A)$ vanishing at 0. More precisely, given an increasing function $\alpha:\Gamma(A)\to[0,\infty)$ with $\alpha(0)=0$, we construct a Choquet trace $\varphi_{\alpha}$ on $A$ which, on each finite-dimensional subalgebra $A_{n}\cong\bigoplus_{s}M_{k_{n,s}}(\mathbb{C})$, admits an explicit Choquet-type formula in terms of the spectrum of $a\in A_{n}^{+}$ and the rank vectors of its spectral projections. Conversely, every Choquet trace on $A$ arises uniquely in this way, and is determined by its values on projections via the associated function $\alpha$ on the scale.
The structure of the paper is as follows. In Section 2, we review the discrete Choquet integral on a finite set and introduce the notion of a Choquet trace on a unital C-algebra. In Section 3, we extend the theory of Nagisa–Watatani on Choquet traces from a single matrix algebra to finite direct sums of matrix algebras. Section 4 is devoted to the AF case: we recall the description of the dimension scale $\Gamma(A)$ of an AF algebra in terms of its inductive limit decomposition, construct Choquet traces from increasing functions on $\Gamma(A)$, and prove the characterization. In Section 5, we illustrate the general theory by discussing concrete examples, including UHF algebras and the Fibonacci AF algebra, where the dimension scale and its order structure can be described explicitly.
2. Choquet integrals and Choquet traces
We begin by recalling the discrete Choquet integral on a finite set, which will serve as a guiding model for the non-commutative constructions considered later. Let $\Omega=\{1,\dots,n\}$ and
$\mu:\mathcal{P}(\Omega)\to[0,\infty)$
be an increasing set function, that is,
$X\subset Y\subset\Omega\ \Longrightarrow\ \mu(X)\leq\mu(Y),\qquad\mu(\emptyset)=0.$
Such a map is usually called a monotone measure or capacity. For a vector $x=(x_{1},\dots,x_{n})\in[0,\infty)^{n}$, choose a permutation $\sigma$ of $\{1,\dots,n\}$ such that
$x_{\sigma(1)}\geq x_{\sigma(2)}\geq\cdots\geq x_{\sigma(n)}.$
Set
$A_{i}:=\{\sigma(1),\dots,\sigma(i)\},\qquad 1\leq i\leq n.$
The discrete Choquet integral of $x$ with respect to $\mu$ is defined by
$(C)\!\int x\,d\mu=\sum_{i=1}^{n-1}\bigl(x_{\sigma(i)}-x_{\sigma(i+1)}\bigr)\,\mu(A_{i})+x_{\sigma(n)}\,\mu(A_{n}).$
This value does not depend on the choice of the permutation $\sigma$ that orders the components in decreasing order.
Real-valued functions $f$ and $g$ on a set $\Omega$ are called comonotonic functions if for all $s,t\in\Omega$
$(f(s)-f(t))(g(s)-g(t))\geq 0.$
The Choquet integral satisfies the following three properties.
  • Monotonicity: if $0\leq f\leq g$, then $(C)\!\int f\,d\mu\leq(C)\!\int g\,d\mu$.
  • Positive homogeneity: if $r\geq 0$, then $(C)\!\int rf\,d\mu=r(C)\!\int f\,d\mu$.
  • Comonotonic additivity: if $f$ and $g$ are comonotonic, then
    $(C)\!\int(f+g)\,d\mu=(C)\!\int f\,d\mu+(C)\!\int g\,d\mu.$
Conversely, every map on $0,\infty)^{n}$ that satisfies these three properties is a Choquet integral with respect to a unique capacity. Thus the Choquet integral is completely characterized by properties of monotonicity, positive homogeneity, and comonotonic additivity.
The characterization above suggests that the Choquet integral can be understood abstractly as a map determined by monotonicity, positive homogeneity, and comonotonic additivity. Our aim is to formulate and study a non-commutative analogue of this notion on positive elements of a C-algebra.
In the non-commutative setting, spectral data of positive elements are naturally reflected in the ordered $K_{0}$-theory. Throughout the paper, for a unital C-algebra AA we write $K_{0}(A)$ for the ordered $K$-group with positive cone $K_{0}(A)^{+}$ and distinguished order unit $1_{A}]\in K_{0}(A)^{+}$.
We define the dimension scale of $A$ by
$\Gamma(A):=\{\,[p]\in K_{0}(A)^{+}:p\in A\text{ is a projection}\,\}.$
Motivated by Nagisa-Watatani’s theory, we now introduce the notion of a Choquet trace on a C-algebra.
Definition 2.1.
Let $A$ be a unital C-algebra. A map $\varphi:A^{+}\to[0,\infty)$ is said to be a Choquet trace if it satisfies the following properties:
  • (U)
    Unitary invariance: for every $a\in A^{+}$ and every unitary $u\in A$,
    $\varphi(uau^{*})=\varphi(a).$
  • (M)
    Monotonicity: if $a,b\in A^{+}$ and $a\leq b$, then $\varphi(a)\leq\varphi(b)$.
  • (H)
    Positive homogeneity: if $r\geq 0$ and $a\in A^{+}$, then
    $\varphi(ra)=r\,\varphi(a).$
  • (C)
    Comonotonic additivity on the spectrum: for each $a\in A^{+}$ and any continuous functions $f,g\in C(\sigma(a))^{+}$ that are comonotonic on $\sigma(a)$,
    $\varphi(f(a)+g(a))=\varphi(f(a))+\varphi(g(a)).$
    Here $f(a)$ and $g(a)$ are defined by continuous functional calculus.
The properties (U), (M), (H), and (C) form an exact non-commutative analogue of the defining properties of the discrete Choquet integral on a finite set. Property (C) expresses comonotonic additivity along the spectrum of a positive element via functional calculus.
The following lemma is proved in [15]. We include a proof for the reader’s convenience.
Lemma 2.2.
Let $A$ be a unital C-algebra and $\varphi:A^{+}\to[0,\infty)$ be a Choquet trace with $\varphi(0)=0$. Then for all $a,b\in A^{+}$,
$|\varphi(a)-\varphi(b)|\leq\varphi(1_{A})\,\|a-b\|.$
In particular, $\varphi$ is norm continuous on $A^{+}$.
Proof.
Let $a,b\in A^{+}$ and put $\varepsilon:=\|a-b\|$. Since aba-b is self-adjoint,
$\varepsilon 1_{A}\leq a-b\leq\varepsilon 1_{A},$
and hence
$a\leq b+\varepsilon 1_{A},\qquad b\leq a+\varepsilon 1_{A}.$
By monotonicity,
$\varphi(a)\leq\varphi(b+\varepsilon 1_{A}),\qquad\varphi(b)\leq\varphi(a+\varepsilon 1_{A}).$
The property (C) implies that
$\varphi(b+\varepsilon 1_{A})=\varphi(b)+\varepsilon\varphi(1_{A}).$
Combining these inequalities yields the claim. ∎
We now recall Nagisa-Watatani’s theory for the matrix algebra $A=M_{n}(\mathbb{C})$ in [11]. Since every projection is Murray–von Neumann equivalent to a diagonal projection of a given rank, we may identify
$\Gamma(A)\;\cong\;\{0,1,\dots,n\},$
where $i\in\{0,\dots,n\}$ corresponds to the $K_{0}$-class of rank-$i$ projections. For notational convenience, we will simply write
$\Gamma(A)=\{0,1,\dots,n\}$
in what follows.
Let
$\mu_{1}(a)\geq\mu_{2}(a)\geq\cdots\geq\mu_{n}(a)$
be the eigenvalues of $a$ in decreasing order with counting multiplicity. Put
$q_{j}(a):=\operatorname{rank}\bigl(E_{a}([\mu_{j}(a),\infty))\bigr)\in\{0,1,\dots,n\},$
where $E_{a}([t,\infty))$ is the spectral projection of $a$ corresponding to the interval $t,\infty)$ for any $t\geq 0$. Define a map $\varphi_{\alpha}:A^{+}\to[0,\infty)$ by
$\varphi_{\alpha}(a)=\sum_{j=1}^{n}\bigl(\mu_{j}(a)-\mu_{j+1}(a)\bigr)\alpha(q_{j}(a)),$
where $\mu_{n+1}(a)=0$. We call $\varphi_{\alpha}$ the Choquet formula associated with $\alpha$. This is defined and characterized in [11], that is, the map $\varphi_{\alpha}$ is a Choquet trace, and conversely every Choquet trace on a matrix algebra can be represented in this form. There is a one-to-one correspondence between Choquet traces on $M_{n}(\mathbb{C})$ and increasing functions on $\Gamma(A)$ that vanish at 0.
3. Non-linear traces on finite direct sums of matrix algebras of Choquet type
In this section, we extend the Nagisa-Watatani’s theory from a single matrix algebra to a finite direct sum
$A\;=\;\bigoplus_{s=1}^{L}M_{k_{s}}(\mathbb{C}).$
The dimension scale $\Gamma(A)$ is identified with the product
$\Gamma(A)\;\cong\;\{0,1,\dots,k_{1}\}\times\cdots\times\{0,1,\dots,k_{L}\}.$
We will henceforth use this identification and simply write
$\Gamma(A)=\{0,1,\dots,k_{1}\}\times\cdots\times\{0,1,\dots,k_{L}\}.$
For a projection $p=\bigoplus_{s=1}^{L}p_{s}\in A$, we define its rank vector by
$\operatorname{rank}(p):=\bigl(\operatorname{rank}(p_{1}),\dots,\operatorname{rank}(p_{L})\bigr)\in\Gamma(A).$
As in Section 2, we define the Choquet trace and the Choquet formula on direct sums of matrix algebras in this section.
Definition 3.1.
For $a\in A$, let
$\lambda_{1}(a)>\lambda_{2}(a)>\cdots>\lambda_{K_{a}}(a)\geq 0$
be the distinct points of the spectrum $\sigma(a)$ listed in decreasing order with $K_{a}=\#\sigma(a)$. Let $\alpha:\Gamma(A)\to[0,\infty)$ be an increasing function such that $\alpha(0)=0$. We define a map $\varphi_{\alpha}:A^{+}\to[0,\infty)$ by
$\varphi_{\alpha}(a)=\sum_{j=1}^{K_{a}}\bigl(\lambda_{j}(a)-\lambda_{j+1}(a)\bigr)\alpha(r_{j}(a)),$
where $\lambda_{K_{a}+1}(a)=0$ and
$r_{i}(a):=[E_{a}([\lambda_{i}(a),\infty))]\in\Gamma(A).$
The Choquet formula in Definition 3.1 expresses $\varphi_{\alpha}(x)$ in terms of the jumps of the spectrum of $x$. For later arguments, especially for the monotonicity of $\varphi_{\alpha}$, it is convenient to allow an arbitrary finite decreasing sequence $t_{1}>\cdots>t_{m}\geq 0$ such that $\sigma(x)\subset\{t_{1},\dots,t_{m}\}$. The following lemma shows that we obtain the same value by replacing the spectral values with such a grid and the corresponding spectral projections $E_{x}([t_{i},\infty))$.
Lemma 3.2.
Let $a\in A^{+}$, and
$t_{1}>t_{2}>\cdots>t_{m}\geq 0$
be any finite decreasing sequence such that
$\sigma(a)\subset\{t_{1},\dots,t_{m}\},$
and put $t_{m+1}:=0$. Then
$\varphi_{\alpha}(a)=\sum_{i=1}^{m}\bigl(t_{i}-t_{i+1}\bigr)\,\alpha\bigl(r(a;t_{i})\bigr),$
where $r(a;t):=\operatorname{rank}\bigl(E_{a}([t,\infty))\bigr)\in\Gamma(A)$.
Proof.
Put $K_{a}:=\#\sigma(a)$, and list the distinct points of $\sigma(a)$ in decreasing order as
$\lambda_{1}(a)>\cdots>\lambda_{K_{a}}(a)\geq 0.$
For each $i$, define
$j(i):=\max\{\,j\in\{1,\dots,K_{a}\}\mid\lambda_{j}(a)\geq t_{i}\,\},$
with the convention that $j(i)=0$ if the set is empty (equivalently, if $t_{i}>\lambda_{1}(a)$). Put $r_{0}(a):=0\in\Gamma(A)$. If $j(i)=0$, then $t_{i}>\lambda_{1}(a)=\|a\|$, and therefore
$E_{a}([t_{i},\infty))=0\quad\text{and}\quad r(a;t_{i})=0=r_{0}(a).$
If $j(i)\geq 1$, then there is no spectral value of $a$ in the open interval $(\lambda_{j(i)+1}(a),\lambda_{j(i)}(a))$, hence
$E_{a}([t_{i},\infty))=E_{a}([\lambda_{j(i)}(a),\infty)),$
and therefore
$r(a;t_{i})=r_{j(i)}(a).$
Hence
$\sum_{i=1}^{m}\bigl(t_{i}-t_{i+1}\bigr)\,\alpha\bigl(r(a;t_{i})\bigr)$ $=\sum_{i=1}^{m}\bigl(t_{i}-t_{i+1}\bigr)\,\alpha\bigl(r_{j(i)}(a)\bigr)$
$=\sum_{j=1}^{K_{a}}\Bigl(\sum_{i:\,j(i)=j}(t_{i}-t_{i+1})\Bigr)\,\alpha\bigl(r_{j}(a)\bigr)$
$=\sum_{j=1}^{K_{a}}\Bigl(\lambda_{j}(a)-\lambda_{j+1}(a)\Bigr)\,\alpha\bigl(r_{j}(a)\bigr)$
$=\varphi_{\alpha}(a)$
The following theorem shows that the map $\varphi_{\alpha}$ defined by Definition 3.1 is a Choquet trace on the finite-dimensional C-algebra $A=\bigoplus_{s=1}^{L}M_{k_{s}}(\mathbb{C})$.
Theorem 3.3.
The map $\varphi_{\alpha}:A^{+}\to[0,\infty)$ is a Choquet trace on $A$.
Proof.
We check the four properties.
(U) For any positive element $a$ and a unitary element $u$ in $A$, we have
$\sigma(a)=\sigma(uau^{*}),\qquad E_{uau^{*}}([t,\infty))=uE_{a}([t,\infty))u^{*}.$
Hence we obtain
$\lambda_{j}(a)=\lambda_{j}(uau^{*}),\qquad r_{j}(uau^{*})=r_{j}(a).$
Therefore the property (U) holds.
(H) Let $t\geq 0$ and $a\in A^{+}$. For any $j$ and $s>0$,
$\lambda_{j}(ta)=t\lambda_{j}(a),\qquad E_{ta}([s,\infty))=E_{a}([s/t,\infty)).$
Thus the rank vectors satisfy
$r_{j}(ta)=r_{j}(a),\qquad\lambda_{j}(ta)-\lambda_{j+1}(ta)=t\bigl(\lambda_{j}(a)-\lambda_{j+1}(a)\bigr),$
and the Choquet formula gives
$\varphi_{\alpha}(ta)=\sum_{j=1}^{K_{a}}\bigl(\lambda_{j}(ta)-\lambda_{j+1}(ta)\bigr)\,\alpha\bigl(r_{j}(ta)\bigr)=t\sum_{j=1}^{K_{a}}\bigl(\lambda_{j}(a)-\lambda_{j+1}(a)\bigr)\,\alpha\bigl(r_{j}(a)\bigr)=t\,\varphi_{\alpha}(a).$
(C) Let $a\in A^{+}$ and $f,g\in C(\sigma(a))^{+}$ be comonotonic on $\sigma(a)$. For simplicity, we denote $\lambda_{j}(a)$ by $\lambda_{j}$ for any $j$. By comonotonicity on the finite totally ordered set $\sigma(a)$, there exists a permutation $\pi$ of $\{1,\dots,K\}$ such that
$f(\lambda_{\pi(1)})\geq\cdots\geq f(\lambda_{\pi(K_{a})})\geq 0,\qquad g(\lambda_{\pi(1)})\geq\cdots\geq g(\lambda_{\pi(K_{a})})\geq 0.$
Put
$f(\lambda_{\pi(K_{a}+1)})=g(\lambda_{\pi(K_{a}+1)})=0.$
Let $h$ be a positive element in $C(\sigma(a))$ such that
$h(\lambda_{\pi(1)}(a))\geq\cdots\geq h(\lambda_{\pi(K_{a})}(a))$
with the convention $h(\lambda_{\pi(K_{a}+1)}(a))=0$. Since we have $\sigma(h(a))=h(\sigma(a))$, we obtain
(1) $\varphi_{\alpha}(h(a))=\sum_{j=1}^{K_{a}}\bigl(h(\lambda_{\pi(j)}(a))-h(\lambda_{\pi(j+1)}(a))\bigr)\alpha\bigl(E_{h(a)}([h(\lambda_{\pi(j)}(a)),\infty))\bigr).$
For any $1\leq i\leq K_{h(a)}$, put
$j_{h}(i):=\max\{j\in\{1,\dots,K_{a}\}\mid\lambda_{i}(h(a))=h(\lambda_{\pi(j)}(a))\},\qquad j_{h}(0)=0.$
Then we have
$E_{h(a)}([h(\lambda_{\pi(j_{h}(i))}(a)),\infty))=E_{h(a)}([\lambda_{i}(h(a)),\infty))=\sum_{j=1}^{j_{h}(i)}E_{a}(\{\lambda_{\pi(j)}(a)\}),$
and
$\sum_{j=1}^{K_{a}}\bigl(h(\lambda_{\pi(j)}(a))-h(\lambda_{\pi(j+1)}(a))\bigr)\beta(j)$
$=\sum_{i=1}^{K_{h(a)}}\bigl(h(\lambda_{j_{h}(i)}(a))-h(\lambda_{j_{h}(i+1)}(a))\bigr)\beta(j_{h}(i))$
for any function $\beta$ on $\{1,\dots,K_{h(a)}\}$.
Using (1), we obtain
$\varphi_{\alpha}((f+g)(a))$ $=\sum_{j=1}^{K_{a}}\bigl((f+g)(\lambda_{\pi(j)}(a))-(f+g)(\lambda_{\pi(j+1)}(a))\bigr)$
$\qquad\times\alpha\!\left(\operatorname{rank}\!\left(E_{(f+g)(a)}([(f+g)(\lambda_{\pi(j)}(a)),\infty))\right)\right)$
$=\sum_{j=1}^{K_{a}}\bigl(f(\lambda_{\pi(j)}(a))-f(\lambda_{\pi(j+1)}(a))\bigr)\alpha\!\left(\operatorname{rank}\!\left(E_{(f+g)(a)}([(f+g)(\lambda_{\pi(j)}(a)),\infty))\right)\right)$
$\quad+\sum_{j=1}^{K_{a}}\bigl(g(\lambda_{\pi(j)}(a))-g(\lambda_{\pi(j+1)}(a))\bigr)\alpha\!\left(\operatorname{rank}\!\left(E_{(f+g)(a)}([(f+g)(\lambda_{\pi(j)}(a)),\infty))\right)\right).$
We claim that for each $j\in\{1,\dots,K_{a}\}$,
$E_{(f+g)(a)}\!\left(\bigl[(f+g)(\lambda_{\pi(j)}(a)),\infty\bigr)\right)=E_{f(a)}\!\left(\bigl[f(\lambda_{\pi(j)}(a)),\infty\bigr)\right).$
Indeed, since $f$ and $g$ are comonotone on $\sigma(a)$ and are ordered by the same permutation $\pi$, we have for every $k$,
$(f+g)(\lambda_{\pi(k)}(a))\geq(f+g)(\lambda_{\pi(j)}(a))\iff f(\lambda_{\pi(k)}(a))\geq f(\lambda_{\pi(j)}(a)).$
Hence,
$\{\lambda\in\sigma(a)\mid(f+g)(\lambda)\geq(f+g)(\lambda_{\pi(j)}(a))\}=\{\lambda\in\sigma(a)\mid f(\lambda)\geq f(\lambda_{\pi(j)}(a))\},$
and by the spectral decomposition $E_{h(a)}([t,\infty))=\sum_{\lambda\in\sigma(a),\,h(\lambda)\geq t}E_{a}(\{\lambda\})$ for $h\in C(\sigma(a))$, we obtain
$E_{(f+g)(a)}\!\left(\bigl[(f+g)(\lambda_{\pi(j)}(a)),\infty\bigr)\right)=E_{f(a)}\!\left(\bigl[f(\lambda_{\pi(j)}(a)),\infty\bigr)\right).$
Thus, considering only the $f$-part, we have
$\sum_{j=1}^{K_{a}}\bigl(f(\lambda_{\pi(j)}(a))-f(\lambda_{\pi(j+1)}(a))\bigr)\alpha\!\left(\operatorname{rank}\!\left(E_{(f+g)(a)}([(f+g)(\lambda_{\pi(j)}(a)),\infty))\right)\right)$
$=\sum_{i=1}^{K_{f(a)}}\bigl(f(\lambda_{\pi(j_{f}(i))}(a))-f(\lambda_{\pi(j_{f}(i)+1)}(a))\bigr)\alpha\!\left(\operatorname{rank}\!\left(E_{(f+g)(a)}([(f+g)(\lambda_{\pi(j_{f}(i))}(a)),\infty))\right)\right)$
$=\sum_{i=1}^{K_{f(a)}}\bigl(f(\lambda_{\pi(j_{f}(i))}(a))-f(\lambda_{\pi(j_{f}(i)+1)}(a))\bigr)\alpha\!\left(\operatorname{rank}\!\left(E_{f(a)}([f(\lambda_{\pi(j_{f}(i))}(a)),\infty))\right)\right)$
$=\sum_{j=1}^{K_{a}}\bigl(f(\lambda_{\pi(j)}(a))-f(\lambda_{\pi(j+1)}(a))\bigr)\alpha\!\left(\operatorname{rank}\!\left(E_{f(a)}([f(\lambda_{\pi(j)}(a)),\infty))\right)\right)$
$=\varphi_{\alpha}(f(a)).$
Similarly,
$\sum_{j=1}^{K_{a}}\bigl(g(\lambda_{\pi(j)}(a))-g(\lambda_{\pi(j+1)}(a))\bigr)\alpha\!\left(\operatorname{rank}\!\left(E_{(f+g)(a)}([(f+g)(\lambda_{\pi(j)}(a)),\infty))\right)\right)$
$=\varphi_{\alpha}(g(a)),$
and hence the map $\phi_{\alpha}$ satisfies
$\varphi_{\alpha}(f(a)+g(a))=\varphi_{\alpha}(f(a))+\varphi_{\alpha}(g(a)).$
(M) Let $0\leq a\leq b$ in $A^{+}$ such that
$a=\bigoplus_{s=1}^{L}a_{s},\qquad b=\bigoplus_{s=1}^{L}b_{s},$
with $a_{s},b_{s}\in M_{k_{s}}(\mathbb{C})^{+}$. Then $0\leq a_{s}\leq b_{s}$ for each $1\leq s\leq L$.
Let
$\mu^{(s)}_{1}(a_{s})\geq\cdots\geq\mu^{(s)}_{k_{s}}(a_{s}),\qquad\mu^{(s)}_{1}(b_{s})\geq\cdots\geq\mu^{(s)}_{k_{s}}(b_{s})$
be the eigenvalues of $a_{s}$ and $b_{s}$, respectively, listed in decreasing order with counting multiplicity. By Weyl’s monotonicity theorem (see [1]), we have
$\mu^{(s)}_{i}(a_{s})\leq\mu^{(s)}_{i}(b_{s})\qquad(1\leq i\leq k_{s},\ 1\leq s\leq L).$
Then, for each $1\leq s\leq L$ and $t\geq 0$,
$\operatorname{rank}\bigl(E_{a_{s}}([t,\infty))\bigr)=\#\{\,i\mid\mu^{(s)}_{i}(a_{s})\geq t\,\}\leq\#\{\,i\mid\mu^{(s)}_{i}(b_{s})\geq t\,\}=\operatorname{rank}\bigl(E_{b_{s}}([t,\infty))\bigr).$
Hence the rank vectors satisfy
$\operatorname{rank}\bigl(E_{a}([t,\infty))\bigr)\leq\operatorname{rank}\bigl(E_{b}([t,\infty))\bigr)\quad\text{in }\Gamma(A)\qquad(t\geq 0).$
Let
$T:=\sigma(a)\cup\sigma(b)\cup\{0\},$
and list its elements in decreasing order as
$t_{1}>t_{2}>\cdots>t_{m}\geq 0,\qquad t_{m+1}:=0.$
Applying Lemma 3.2 to $a$ and $b$ with this sequence, and using the monotonicity of $\alpha$ on $\Gamma(A)$, we obtain
$\varphi_{\alpha}(a)$ $=\sum_{i=1}^{m}(t_{i}-t_{i+1})\,\alpha\bigl(\operatorname{rank}(E_{a}([t_{i},\infty)))\bigr)$
$\leq\sum_{i=1}^{m}(t_{i}-t_{i+1})\,\alpha\bigl(\operatorname{rank}(E_{b}([t_{i},\infty)))\bigr)=\varphi_{\alpha}(b).$
Thus $\varphi_{\alpha}$ is monotone. ∎
We now show that the above construction exhausts all Choquet traces on the finite-dimensional C-algebra $A=\bigoplus_{s=1}^{L}M_{k_{s}}(\mathbb{C})$.
Theorem 3.4.
Let $A=\bigoplus_{s=1}^{L}M_{K_{s}}(\mathbb{C})$, and $\varphi:A^{+}\to[0,\infty)$ be a Choquet trace on $A$. Then there exists a unique increasing function $\alpha:\Gamma(A)\to[0,\infty)$ with $\alpha(0)=0$ such that
$\varphi(a)=\varphi_{\alpha}(a)\qquad(a\in A^{+}).$
Proof.
Assume that $\varphi:A^{+}\to[0,\infty)$ satisfies (U), (M), (H) and (C).
For $x\in\Gamma(A)$, let $p(x)\in A$ be a projection satisfying $\operatorname{rank}(p(x))=x$, and define
$\alpha(x):=\varphi(p(x)).$
We first check that $\alpha$ is well-defined. If $p,q\in A$ are projections with $\operatorname{rank}(p)=\operatorname{rank}(q)=x$, then $p$ and $q$ are unitarily equivalent in $A$. By unitary invariance (U), we have $\varphi(p)=\varphi(q)$. Hence $\alpha$ is well-defined. In particular, $\alpha(0)=\varphi(0)=0$.
Next we show that $\alpha$ is increasing. Let $x,y\in\Gamma(A)$ with $x\leq y$. Choose projections $p,q\in A$ such that $\operatorname{rank}(p)=x$, $\operatorname{rank}(q)=y$, and $p\leq q$. By monotonicity (M),
$\alpha(x)=\varphi(p)\leq\varphi(q)=\alpha(y).$
Thus $\alpha:\Gamma(A)\to[0,\infty)$ is increasing.
Let $a\in A^{+}$. To simplify notation, we denote $\lambda_{j}(a)$ by $\lambda_{j}$ for all $j$, and set $\lambda_{K_{a}+1}:=0$. For $x\in\sigma(a)$ and $1\leq k\leq K_{a}$, define
$s_{k}(x):=(\lambda_{k}-\lambda_{k+1})\,\mathbf{1}_{[\lambda_{k},\infty)}(x).$
Then each $s_{k}$ is monotone increasing on $\sigma(a)$.
On $\sigma(a)$, we have the identity
$x=\sum_{k=1}^{K_{a}}s_{k}(x)\qquad(x\in\sigma(a)).$
By continuous functional calculus,
$a=\sum_{k=1}^{K_{a}}s_{k}(a).$
We claim that
$\varphi\Bigl(\sum_{k=1}^{K_{a}}s_{k}(a)\Bigr)=\sum_{k=1}^{K_{a}}\varphi\bigl(s_{k}(a)\bigr).$
Indeed, set $b_{1}:=s_{1}(a)$ and $b_{m}:=b_{m-1}+s_{m}(a)$ for $m\geq 2$. Since $s_{m}$ and $s_{1}+\cdots+s_{m-1}$ are both monotone increasing on $\sigma(a)$, they are comonotonic. Hence by (C),
$\varphi(b_{m})=\varphi(b_{m-1})+\varphi\bigl(s_{m}(a)\bigr).$
By induction on $m$, we obtain the claim.
Using (H), we have
$\varphi\bigl(s_{k}(a)\bigr)=(\lambda_{k}-\lambda_{k+1})\,\varphi\bigl(\mathbf{1}_{[\lambda_{k},\infty)}(a)\bigr).$
Since $\mathbf{1}_{[\lambda_{k},\infty)}(a)=E_{a}([\lambda_{k},\infty))$,
$\varphi\bigl(\mathbf{1}_{[\lambda_{k},\infty)}(a)\bigr)=\alpha\bigl(\operatorname{rank}(E_{a}([\lambda_{k},\infty)))\bigr)=\alpha\bigl(r_{k}(a)\bigr).$
Therefore,
$\varphi(a)=\sum_{k=1}^{K_{a}}(\lambda_{k}-\lambda_{k+1})\,\alpha\bigl(r_{k}(a)\bigr)=\varphi_{\alpha}(a).$
4. Non-linear traces on AF algebras of Choquet type
In this section, we extend the Choquet formula from finite-dimensional C-algebras to general AF algebras. Throughout this section, an AF algebra means a unital AF algebra.
Let $A$ be a unital AF algebra. There exists an increasing sequence of finite-dimensional C-subalgebras
$A_{1}\subset A_{2}\subset\cdots\subset A_{n}\subset\cdots\subset A,\qquad A=\overline{\bigcup_{n\geq 1}A_{n}},$
where each $A_{n}$ is of the form
$A_{n}=\bigoplus_{s=1}^{L_{n}}M_{k_{n,s}}(\mathbb{C}).$
Since $A$ is unital, we may and do assume that the inductive system $(A_{n})_{n\geq 1}$ consists of unital embeddings, so that the unit of each $A_{n}$ is identified with the unit $1_{A}$ of $A$.
For $m\geq n$, we denote by
$\rho_{n,m}:A_{n}\longrightarrow A_{m}$
the canonical connecting *-homomorphism (the inclusion map), and we set
$\rho_{n,n}:=\mathrm{id}_{A_{n}}.$
Then $\rho_{n,\ell}=\rho_{m,\ell}\circ\rho_{n,m}$ holds for $\ell\geq m\geq n$.
Each $\rho_{n,m}$ induces an order-preserving map on the dimension scales
$\gamma_{n,m}:\Gamma(A_{n})\longrightarrow\Gamma(A_{m}),\qquad\gamma_{n,m}([p]):=[\rho_{n,m}(p)]$
for a projection $p\in A_{n}$. Then we have
$\gamma_{n,n}:=\mathrm{id}_{\Gamma(A_{n})}.$
Then $\gamma_{n,\ell}=\gamma_{m,\ell}\circ\gamma_{n,m}$ holds for $\ell\geq m\geq n$.
It is well known that the ordered group $K_{0}(A)$ is the inductive limit of $(K_{0}(A_{n}),(\rho_{n,m})_{*})$ and that the dimension scale $\Gamma(A)$ is the inductive limit
[…]
in the category of ordered sets (see [3, 6]). We denote by
$\gamma_{n}:\Gamma(A_{n})\longrightarrow\Gamma(A)$
the canonical map into the inductive limit. By construction,
$\gamma_{\ell}\circ\gamma_{n,\ell}=\gamma_{n}\qquad(\ell\geq n).$
Moreover, since the embeddings are unital, the element $\gamma_{n}([1_{A_{n}}])\in\Gamma(A)$ does not depend on $n$ and coincides with the order unit $1_{A}]$.
In the finite-dimensional case, Choquet traces are completely described by increasing functions on the corresponding finite dimension scales. This observation provides the guiding principle for the AF case.
Let $\alpha:\Gamma(A)\to[0,\infty)$ be an increasing map with $\alpha(0)=0$. For each $n$, define
$\alpha^{(n)}:\Gamma(A_{n})\longrightarrow[0,\infty),\qquad\alpha^{(n)}(x):=\alpha(\gamma_{n}(x)).$
By section 3, there exists a unique Choquet trace
$\varphi_{\alpha}^{(n)}:A_{n}^{+}\longrightarrow[0,\infty)$
such that
$\varphi_{\alpha}^{(n)}(p)=\alpha^{(n)}(\operatorname{rank}(p))$
for any projection $p\in A_{n}$. The Choquet formula takes the form
$\varphi_{\alpha}^{(n)}(a)=\sum_{k=1}^{K_{a}}(\lambda_{k}(a)-\lambda_{k+1}(a))\,\alpha^{(n)}(r_{k}(a)),$
where $r_{k}(a)=\operatorname{rank}(E_{a}([\lambda_{k}(a),\infty)))\in\Gamma(A_{n})$ and $\lambda_{K_{a}+1}(a):=0$.
In order to pass from the finite-dimensional algebras $A_{n}$ to the AF algebra $A$, we need to verify that the Choquet traces constructed at each stage are compatible with the connecting maps. The following lemma establishes this compatibility.
Lemma 4.1.
For every $n$ and every $a\in A_{n}^{+}$,
$\varphi_{\alpha}^{(n)}(a)=\varphi_{\alpha}^{(n+1)}(\rho_{n,n+1}(a)).$
Proof.
Let $a\in A_{n}^{+}$. List the distinct spectral values of $a$ in decreasing order as
$\lambda_{1}(a)>\cdots>\lambda_{K_{a}}(a)\geq 0,$
and put $\lambda_{K_{a}+1}(a):=0$. Since $\rho_{n,n+1}$ is a *-homomorphism, the functional calculus implies
$\sigma(\rho_{n,n+1}(a))=\sigma(a),$
and hence we may write
$\lambda_{j}(\rho_{n,n+1}(a))=\lambda_{j}(a)\quad(j=1,\dots,K_{a}).$
For each $j$, the functional calculus gives
$E_{\rho_{n,n+1}(a)}([\lambda_{j}(a),\infty))=\rho_{n,n+1}(E_{a}([\lambda_{j}(a),\infty))).$
Taking ranks and using the definition of $\gamma_{n,n+1}$, we obtain
$\operatorname{rank}\!\left(E_{\rho_{n,n+1}(a)}([\lambda_{j}(a),\infty))\right)=\gamma_{n,n+1}\!\left(\operatorname{rank}(E_{a}([\lambda_{j}(a),\infty)))\right).$
Applying $\gamma_{n+1}$ and using $\gamma_{n+1}\circ\gamma_{n,n+1}=\gamma_{n}$, we have
$\gamma_{n+1}\!\left(\operatorname{rank}(E_{\rho_{n,n+1}(a)}([\lambda_{j}(a),\infty)))\right)=\gamma_{n}\!\left(\operatorname{rank}(E_{a}([\lambda_{j}(a),\infty)))\right).$
Therefore,
$\varphi_{\alpha}^{(n+1)}(\rho_{n,n+1}(a))$ $=\sum_{j=1}^{K_{a}}(\lambda_{j}(a)-\lambda_{j+1}(a))\,\alpha\!\left(\gamma_{n+1}\!\left(\operatorname{rank}(E_{\rho_{n,n+1}(a)}([\lambda_{j}(a),\infty)))\right)\right)$
$=\sum_{j=1}^{K_{a}}(\lambda_{j}(a)-\lambda_{j+1}(a))\,\alpha\!\left(\gamma_{n}\!\left(\operatorname{rank}(E_{a}([\lambda_{j}(a),\infty)))\right)\right)$
$=\varphi_{\alpha}^{(n)}(a),$
which proves the assertion. ∎
To extend the local constructions to the whole algebra, we need a spectral approximation lemma ensuring that spectra are preserved in the approximation process.
Lemma 4.2.
Let $A$ be a unital AF algebra with an increasing sequence $(A_{n})_{n\geq 1}$ of finite-dimensional subalgebras whose union is dense in $A$. For any $a\in A^{+}$, there exists a sequence $(a_{m})_{m\geq 1}\subset\bigcup_{n}A_{n}^{+}$ such that
$\|a_{m}\|\leq\|a\|,\qquad\|a_{m}-a\|\to 0,\qquad\sigma(a_{m})\subset\sigma(a).$
Proof.
For each $m\in\mathbb{N}$, choose $n(m)$ and $b_{m}\in A_{n(m)}^{+}$ such that
$\|a-b_{m}\|<\frac{1}{m},\qquad\|b_{m}\|\leq\|a\|.$
Since $b_{m}$ is positive and finite-dimensional, it admits a spectral decomposition
$b_{m}=\sum_{i=1}^{r_{m}}\lambda_{i}^{(m)}p_{i}^{(m)},$
where $p_{i}^{(m)}\in A_{n(m)}$ are pairwise orthogonal projections and $\lambda_{i}^{(m)}\geq 0$.
For each $i$, choose $\mu_{i}^{(m)}\in\sigma(a)$ such that
$|\mu_{i}^{(m)}-\lambda_{i}^{(m)}|=dist(\lambda_{i}^{(m)},\sigma(a)).$
Define
$a_{m}:=\sum_{i=1}^{r_{m}}\mu_{i}^{(m)}p_{i}^{(m)}\in A_{n(m)}^{+}.$
Then $\sigma(a_{m})\subset\sigma(a)$ by construction. Moreover,
$\|a_{m}-b_{m}\|=\max_{i}|\mu_{i}^{(m)}-\lambda_{i}^{(m)}|\leq\|a-b_{m}\|<\frac{1}{m},$
and hence
$\|a_{m}-a\|\leq\|a_{m}-b_{m}\|+\|b_{m}-a\|<\frac{2}{m}.$
Finally, since $\mu_{i}^{(m)}\leq\|a\|$ for all $i$, we have $\|a_{m}\|\leq\|a\|$. This completes the proof. ∎
To extend the locally defined maps to the whole algebra $A$, we need a uniform continuity estimate. We obtain a uniform Lipschitz bound for the local maps $\varphi_{\alpha}^{(n)}$ from Lemma 2.2. In particular, the map
$\bigcup_{n\geq 1}A_{n}^{+}\longrightarrow[0,\infty),\qquad a\longmapsto\varphi_{\alpha}^{(n)}(a)\ (a\in A_{n}^{+}),$
is well defined and $L$-Lipschitz.
We now construct a global Choquet trace on $A$ from an increasing function on the dimension scale.
Theorem 4.3.
Let $A$ be an AF algebra and $\alpha:\Gamma(A)\to[0,\infty)$ be an increasing map with $\alpha(0)=0$. Then there exists a unique Choquet trace
$\Phi_{\alpha}:A^{+}\longrightarrow[0,\infty)$
such that
$\Phi_{\alpha}(p)=\alpha([p])\qquad\text{for every projection }p\in A.$
Proof.
Define a map
$f:\bigcup_{n\geq 1}A_{n}^{+}\to[0,\infty),\qquad f(a):=\varphi_{\alpha}^{(n)}(a)\quad(a\in A_{n}^{+}).$
By Lemma 4.1, $f$ is well defined. Moreover, each $\varphi_{\alpha}^{(n)}$ is $L$-Lipschitz with
$L=\alpha([1_{A}]),$
hence $f$ is $L$-Lipschitz on $\bigcup_{n}A_{n}^{+}$. Since $\bigcup_{n}A_{n}^{+}$ is norm-dense in $A^{+}$, $f$ extends uniquely to a $L$-Lipschitz map
$\Phi_{\alpha}:A^{+}\to[0,\infty).$
Clearly, $\Phi_{\alpha}(p)=\alpha([p])$ holds for every projection $p\in A$.
(U) Let $u\in A$ be unitary and $a\in A^{+}$. Choose a sequence $n(k)$ and elements $u_{k}\in A_{n(k)}$ and $a_{k}\in A_{n(k)}^{+}$ such that $\|u-u_{k}\|\to 0$ and $\|a-a_{k}\|\to 0$. For sufficiently large $k$, replace $u_{k}$ by its polar part $u_{k}(u_{k}^{*}u_{k})^{-1/2}$, so that $u_{k}$ is unitary and still $u_{k}\to u$ in norm. Since $\varphi_{\alpha}^{(n(k))}$ is unitarily invariant on $A_{n(k)}^{+}$,
$\Phi_{\alpha}(a_{k})=\varphi_{\alpha}^{(n(k))}(a_{k})=\varphi_{\alpha}^{(n(k))}(u_{k}a_{k}u_{k}^{*})=\Phi_{\alpha}(u_{k}a_{k}u_{k}^{*}).$
Letting $k\to\infty$ and using norm-continuity of $\Phi_{\alpha}$, we obtain $\Phi_{\alpha}(a)=\Phi_{\alpha}(uau^{*})$.
(M) Let $0\leq a\leq b$ in $A$. Choose $n(k)$ and $a_{k},b_{k}\in A_{n(k)}^{+}$ such that $\|a-a_{k}\|\to 0$ and $\|b-b_{k}\|\to 0$. Fix $\varepsilon>0$. For sufficiently large $k$, we have $\|a-a_{k}\|<\varepsilon$ and $\|b-b_{k}\|<\varepsilon$. Since $a\leq b$, it follows that
$a_{k}\leq b_{k}+2\varepsilon 1_{A}$
for sufficiently large $k$. Using monotonicity and positive homogeneity on $A_{n(k)}^{+}$,
$\Phi_{\alpha}(a_{k})=\varphi_{\alpha}^{(n(k))}(a_{k})\leq\varphi_{\alpha}^{(n(k))}(b_{k}+2\varepsilon 1_{A})=\Phi_{\alpha}(b_{k}+2\varepsilon 1_{A}).$
Letting $k\to\infty$ and using continuity, we obtain
$\Phi_{\alpha}(a)\leq\Phi_{\alpha}(b+2\varepsilon 1_{A}),$
and hence we have $\Phi_{\alpha}(a)\leq\Phi_{\alpha}(b)$.
(H) Let $r\geq 0$ and $a\in A^{+}$. Choose $a_{k}\in\bigcup_{n}A_{n}^{+}$ with $\|a-a_{k}\|\to 0$. Then for each $k$,
$\Phi_{\alpha}(ra_{k})=\varphi_{\alpha}^{(n(k))}(ra_{k})=r\,\varphi_{\alpha}^{(n(k))}(a_{k})=r\,\Phi_{\alpha}(a_{k}).$
Letting $k\to\infty$, we obtain $\Phi_{\alpha}(ra)=r\,\Phi_{\alpha}(a)$.
(C) Let $a\in A^{+}$ and any continuous functions $f,g\in C(\sigma(a))^{+}$ that are comonotonic on $\sigma(a)$. Choose a sequence $a_{m}\in\bigcup_{n}A_{n}^{+}$ such that
$\|a_{m}-a\|\to 0,\qquad\sigma(a_{m})\subset\sigma(a)$
(Lemma 4.2). Then $f$ and $g$ are also comonotone on $\sigma(a_{m})$.
For $m$, let $n(m)$ with $a_{m}\in A_{n(m)}^{+}$. Since the property (C) of $\varphi_{\alpha}^{(n(m))}$, we have
$\Phi_{\alpha}\bigl(f(a_{m})+g(a_{m})\bigr)$ $=\varphi_{\alpha}^{(n(m))}\bigl(f(a_{m})+g(a_{m})\bigr)$
$=\varphi_{\alpha}^{(n(m))}\bigl(f(a_{m})\bigr)+\varphi_{\alpha}^{(n(m))}\bigl(g(a_{m})\bigr)$
$=\Phi_{\alpha}\bigl(f(a_{m})\bigr)+\Phi_{\alpha}\bigl(g(a_{m})\bigr).$
Since $f$ and $g$ are continuous on $0,\|a\|]$, we obtain $f(a_{m})\to f(a)$ and $g(a_{m})\to g(a)$ in norm. By continuity of $\Phi_{\alpha}$, letting $m\to\infty$ yields
$\Phi_{\alpha}\bigl(f(a)+g(a)\bigr)=\Phi_{\alpha}\bigl(f(a)\bigr)+\Phi_{\alpha}\bigl(g(a)\bigr).$
Therefore $\Phi_{\alpha}$ is a Choquet trace.
Finally, we show that the map $\Phi_{\alpha}$ is unique. Let $\Psi:A^{+}\to[0,\infty)$ be a Choquet trace with $\Psi(p)=\alpha([p])$ for every projection $p\in A$. Fix $n$, the characterization of Choquet traces on $A_{n}$ shows that $\Psi|_{A_{n}^{+}}=\varphi_{\alpha}^{(n)}$. Hence $\Psi=f$ on $\bigcup_{n}A_{n}^{+}$, and by continuity, we conclude $\Psi=\Phi_{\alpha}$ on $A^{+}$. ∎
We next show that every Choquet trace on $A$ arises uniquely from an increasing function on the dimension scale.
Lemma 4.4.
Let $\varphi:A^{+}\to[0,\infty)$ be a Choquet trace on an AF algebra $A$. Then there exists a unique increasing map
$\alpha:\Gamma(A)\to[0,\infty),\qquad\alpha(0)=0,$
such that
$\alpha([p])=\varphi(p)\qquad\text{for every projection }p\in A.$
Proof.
Define $\alpha$ on classes of projections by
$\alpha([p]):=\varphi(p)\qquad(p\ \text{a projection in }A).$
We first show that this is well defined.
Let $p,q$ be projections in $A$ with $[p]=[q]$ in $\Gamma(A)$. By the fact that $\Gamma(A)$ (equivalently $K_{0}(A)^{+}$) is the inductive limit of $\Gamma(A_{n})$, there exist $n$ and projections $p^{\prime},q^{\prime}\in A_{n}$ such that
$p]=\gamma_{n}([p^{\prime}]),\qquad[q]=\gamma_{n}([q^{\prime}])\quad\text{in }\Gamma(A),$
and
$p^{\prime}]=[q^{\prime}]\quad\text{in }\Gamma(A_{n}).$
In the finite-dimensional algebra $A_{n}$, equality of classes implies that $p^{\prime}$ and $q^{\prime}$ are Murray–von Neumann equivalent, hence unitarily equivalent. Since $\varphi$ is unitarily invariant, we obtain $\varphi(p^{\prime})=\varphi(q^{\prime})$.
On the other hand, $p]=\gamma_{n}([p^{\prime}])$ means that $p$ is Murray–von Neumann equivalent to the image of $p^{\prime}$ in $A$ (and similarly for $q$ and $q^{\prime}$); in particular, $\varphi(p)=\varphi(p^{\prime})$ and $\varphi(q)=\varphi(q^{\prime})$ by unitary invariance. Therefore $\varphi(p)=\varphi(q)$, proving that $\alpha$ is well defined.
Clearly $\alpha(0)=0$. To see that $\alpha$ is increasing, suppose $p]\leq[q]$ in $\Gamma(A)$. Then at some stage $A_{n}$ we can represent these classes by projections $p^{\prime},q^{\prime}\in A_{n}$ with $p^{\prime}]\leq[q^{\prime}]$ in $\Gamma(A_{n})$. Thus $p^{\prime}$ is Murray–von Neumann subequivalent to $q^{\prime}$ in $A_{n}$, and hence $p^{\prime}\precsim q^{\prime}$ implies $\varphi(p^{\prime})\leq\varphi(q^{\prime})$ by monotonicity of $\varphi$. Consequently, $\alpha([p])\leq\alpha([q])$.
Finally, uniqueness is immediate from the defining relation $\alpha([p])=\varphi(p)$ for projections $p$. ∎
Theorem 4.5.
There is a bijection between the set of increasing maps on $\Gamma(A)$ and the set of Choquet traces on A, under which each increasing map $\alpha$ corresponds to the Choquet trace $\varphi_{\alpha}$.
Proof.
Let
$\mathcal{A}:=\{\alpha:\Gamma(A)\to[0,\infty)\mid\alpha\text{ is increasing and }\alpha(0)=0\}$
and
$\mathcal{C}:=\{\text{Choquet traces on }A\}.$
Define a map
$F:\mathcal{A}\to\mathcal{C},\qquad F(\alpha):=\Phi_{\alpha},$
where $\Phi_{\alpha}$ is the Choquet trace constructed in Theorem 4.3.
Suppose $\alpha_{1},\alpha_{2}\in\mathcal{A}$ satisfy $F(\alpha_{1})=F(\alpha_{2})$. Let $p\in A$ be any projection. Then
$\alpha_{1}([p])=\Phi_{\alpha_{1}}(p)=\Phi_{\alpha_{2}}(p)=\alpha_{2}([p]).$
Since every element of $\Gamma(A)$ is represented by the class of a projection, it follows that $\alpha_{1}=\alpha_{2}$. Hence $F$ is injective.
Let $\varphi\in\mathcal{C}$ be an arbitrary Choquet trace on $A$. By Lemma 4.4, there exists an increasing map $\alpha\in\mathcal{A}$ such that
$\alpha([p])=\varphi(p)\qquad\text{for every projection }p\in A.$
We claim that $\varphi=\Phi_{\alpha}$.
The restriction $\varphi|_{A_{n}^{+}}$ is a Choquet trace on the finite-dimensional $C^{*}$-algebra $A_{n}$ for any $n\geq 1$. By the finite-dimensional characterization of Choquet traces,
$\varphi|_{A_{n}^{+}}=\varphi_{\alpha}^{(n)}.$
Since this holds for every $n$, the two maps $\varphi$ and $\Phi_{\alpha}$ coincide on $\bigcup_{n\geq 1}A_{n}^{+}$. As $\bigcup_{n\geq 1}A_{n}^{+}$ is norm-dense in $A^{+}$ and both maps are norm-continuous, we conclude that
$\varphi=\Phi_{\alpha}\quad\text{on }A^{+}.$
Thus $F$ is surjective. ∎
5. Examples
We illustrate Theorem 4.5 by several concrete examples of dimension scales arising from AF algebras. In each case, Choquet traces are described in terms of increasing maps on the corresponding scale. For the computations of the $K_{0}$-groups and the associated dimension scales, we refer to [6], [3].
Example 5.1 (UHF algebras).
Let $A$ be a UHF algebra of type
$N=\prod_{p}p^{n_{p}},\qquad n_{p}\in\mathbb{N}\cup\{\infty\}.$
It is well known that $A$ can be realized as the inductive limit of full matrix algebras
$A_{k}\cong M_{N_{k}}(\mathbb{C}),\qquad N_{k}=n_{1}\cdots n_{k},$
with unital connecting maps. At each finite stage $k$, the dimension scale is given by
$\Gamma(A_{k})=\{0,1,\dots,N_{k}\}.$
Let $\mathbb{Q}_{N}$ denote the subgroup of $\mathbb{Q}$ consisting of rational numbers whose denominators divide $N$. In the inductive limit, the dimension scale $\Gamma(A)$ can be identified, as an ordered subset of $\mathbb{R}$, with
$\Gamma(A)\cong[0,1]\cap\mathbb{Q}_{N},$
via the correspondence
$\gamma_{k}(\operatorname{rank}(p))\longmapsto\frac{\operatorname{rank}(p)}{N_{k}},\qquad p\in A_{k},$
where the order unit corresponds to $1$.
The algebra $A$ admits a unique tracial state $\tau$. By Theorem 4.5, Choquet traces on $A$ are in bijection with increasing maps on $\Gamma(A)$. In particular, any increasing function
$f:[0,1]\to[0,\infty),\qquad f(0)=0,$
induces an increasing map on the scale via
$\alpha_{f}([p]):=f(\tau(p)),$
and hence a Choquet trace $\Phi_{\alpha_{f}}$ on $A$. In general, unless $f$ is linear, we have
$\Phi_{\alpha_{f}}(a)\neq f(\tau(a))$
for $a\in A^{+}$.
Example 5.2 (The Fibonacci AF algebra).
Let $A$ be the Fibonacci AF algebra, whose Bratteli diagram has incidence matrix
$F=\begin{pmatrix}1&1\\ 1&0\end{pmatrix}$
[9]. The ordered $K_{0}$-group of $A$ is given by the inductive limit
[…]
and the dimension scale $\Gamma(A)$ consists of the classes of projections in $K_{0}(A)^{+}$.
The algebra $A$ is simple and admits a unique tracial state $\tau$. The dimension scale $\Gamma(A)$ can be realized as a dense ordered subset of $[0,1]$.
As in the UHF case, Choquet traces on $A$ correspond to increasing maps on $\Gamma(A)$. Although both examples admit a unique tracial state, their dimension scales admit different realizations as ordered subsets of $\mathbb{R}$, which give rise to different classes of increasing maps and hence to different families of Choquet traces.
Example 5.3 (A construction from a tracial state).
Let $A$ be a unital AF algebra admitting a tracial state $\tau$. For $0<\lambda\leq 1$, define an increasing map on $\Gamma(A)$
$\alpha_{\lambda}([p]):=\tau(p)^{\lambda},$
for any projection $p$. By Theorem 4.5, this yields a Choquet trace $\Phi_{\alpha_{\lambda}}$ on $A$. When $\lambda\neq 1$, the resulting Choquet trace is non-linear and is not determined solely by the value of $\tau(a)$ for general $a\in A^{+}$.
These examples illustrate how the order-theoretic structure of the dimension scale $\Gamma(A)$ governs the possible classes of Choquet traces on AF algebras.

### v2

@@
 $\varphi(a)\leq\varphi(b+\varepsilon 1_{A}),\qquad\varphi(b)\leq\varphi(a+\varepsilon 1_{A}).$
-The property (C) implies that
+The properties (C) and (H) imply that
 $\varphi(b+\varepsilon 1_{A})=\varphi(b)+\varepsilon\varphi(1_{A}).$
@@
 There is a one-to-one correspondence between Choquet traces on $M_{n}(\mathbb{C})$ and increasing functions on $\Gamma(A)$ that vanish at 0.
-3. Non-linear traces on finite direct sums of matrix algebras of Choquet type
+3. Non-linear traces of Choquet type on finite direct sums of matrix algebras
 In this section, we extend the Nagisa-Watatani’s theory from a single matrix algebra to a finite direct sum
@@
 For simplicity, we denote $\lambda_{j}(a)$ by $\lambda_{j}$ for any $j$.
-By comonotonicity on the finite totally ordered set $\sigma(a)$, there exists a permutation $\pi$ of $\{1,\dots,K\}$ such that
+By comonotonicity on the finite set $\sigma(a)$, there exists a permutation $\pi$ of $\{1,\dots,K\}$ such that
 $f(\lambda_{\pi(1)})\geq\cdots\geq f(\lambda_{\pi(K_{a})})\geq 0,\qquad g(\lambda_{\pi(1)})\geq\cdots\geq g(\lambda_{\pi(K_{a})})\geq 0.$
@@
 $=\varphi_{\alpha}(g(a)),$
-and hence the map $\phi_{\alpha}$ satisfies
+and hence the map $\varphi_{\alpha}$ satisfies
 $\varphi_{\alpha}(f(a)+g(a))=\varphi_{\alpha}(f(a))+\varphi_{\alpha}(g(a)).$
@@
 $\varphi(a)=\sum_{k=1}^{K_{a}}(\lambda_{k}-\lambda_{k+1})\,\alpha\bigl(r_{k}(a)\bigr)=\varphi_{\alpha}(a).$
-4. Non-linear traces on AF algebras of Choquet type
+4. Non-linear traces of Choquet type on AF algebras
 In this section, we extend the Choquet formula from finite-dimensional C-algebras to general AF algebras. Throughout this section, an AF algebra means a unital AF algebra.
@@
 $\lambda_{1}(a)>\cdots>\lambda_{K_{a}}(a)\geq 0,$
-and put $\lambda_{K_{a}+1}(a):=0$. Since $\rho_{n,n+1}$ is a *-homomorphism, the functional calculus implies
+and put $\lambda_{K_{a}+1}(a):=0$. Since $\rho_{n,n+1}$ is a unital *-homomorphism, we have
 $\sigma(\rho_{n,n+1}(a))=\sigma(a),$
@@
 Letting $k\to\infty$ and using continuity, we obtain
-$\Phi_{\alpha}(a)\leq\Phi_{\alpha}(b+2\varepsilon 1_{A}),$
-and hence we have $\Phi_{\alpha}(a)\leq\Phi_{\alpha}(b)$.
+$\Phi_{\alpha}(a)\leq\Phi_{\alpha}(b+2\varepsilon 1_{A}).$
+Since $\varepsilon$ is arbitrary, we have $\Phi_{\alpha}(a)\leq\Phi_{\alpha}(b)$.
 (H) Let $r\geq 0$ and $a\in A^{+}$. Choose $a_{k}\in\bigcup_{n}A_{n}^{+}$ with $\|a-a_{k}\|\to 0$. Then for each $k$,
@@
 $=\Phi_{\alpha}\bigl(f(a_{m})\bigr)+\Phi_{\alpha}\bigl(g(a_{m})\bigr).$
-Since $f$ and $g$ are continuous on $0,\|a\|]$, we obtain $f(a_{m})\to f(a)$ and $g(a_{m})\to g(a)$ in norm. By continuity of $\Phi_{\alpha}$, letting $m\to\infty$ yields
+Since $f$ and $g$ are continuous on $\sigma(a)$, we obtain $f(a_{m})\to f(a)$ and $g(a_{m})\to g(a)$ in norm. By continuity of $\Phi_{\alpha}$, letting $m\to\infty$ yields
 $\Phi_{\alpha}\bigl(f(a)+g(a)\bigr)=\Phi_{\alpha}\bigl(f(a)\bigr)+\Phi_{\alpha}\bigl(g(a)\bigr).$
@@
 Let $\Psi:A^{+}\to[0,\infty)$ be a Choquet trace with $\Psi(p)=\alpha([p])$ for every projection $p\in A$.
+By Lemma 2.2, $\Psi$ is Lipschitz continuous.
 Fix $n$, the characterization of Choquet traces on $A_{n}$ shows that $\Psi|_{A_{n}^{+}}=\varphi_{\alpha}^{(n)}$.
@@
 Proof.
-Define $\alpha$ on classes of projections by
-$\alpha([p]):=\varphi(p)\qquad(p\ \text{a projection in }A).$
+Define $\alpha$ on $\Gamma(A)$ by
+$\alpha([p]):=\varphi(p)\qquad\text{for every }[p]\in\Gamma(A).$
 We first show that this is well defined.
-Let $p,q$ be projections in $A$ with $[p]=[q]$ in $\Gamma(A)$.
-By the fact that $\Gamma(A)$ (equivalently $K_{0}(A)^{+}$) is the inductive limit of $\Gamma(A_{n})$, there exist $n$ and projections $p^{\prime},q^{\prime}\in A_{n}$ such that
-$p]=\gamma_{n}([p^{\prime}]),\qquad[q]=\gamma_{n}([q^{\prime}])\quad\text{in }\Gamma(A),$
-and
-$p^{\prime}]=[q^{\prime}]\quad\text{in }\Gamma(A_{n}).$
-In the finite-dimensional algebra $A_{n}$, equality of classes implies that $p^{\prime}$ and $q^{\prime}$ are Murray–von Neumann equivalent, hence unitarily equivalent.
-Since $\varphi$ is unitarily invariant, we obtain $\varphi(p^{\prime})=\varphi(q^{\prime})$.
-On the other hand, $p]=\gamma_{n}([p^{\prime}])$ means that $p$ is Murray–von Neumann equivalent to the image of $p^{\prime}$ in $A$ (and similarly for $q$ and $q^{\prime}$); in particular, $\varphi(p)=\varphi(p^{\prime})$ and $\varphi(q)=\varphi(q^{\prime})$ by unitary invariance.
-Therefore $\varphi(p)=\varphi(q)$, proving that $\alpha$ is well defined.
+Let $p,q\in A$ be Murray–von Neumann equivalent projections. Since $A$ is an AF algebra, we have
+$1-p]=[1-q]\quad\text{in }K_{0}(A).$
+Hence $p$ and $q$ are unitarily equivalent. Therefore, the map $\alpha$ is well-defined.
 Clearly $\alpha(0)=0$.